# Discrete-state Continuous-time Diffusion for Graph Generation

### v1

@@ 4.2 Molecule Graph Generation
DisCo-MPNN 83.9 100.0 98.8 87.3 1.63 0.48 13.5
Table 4: Generation performance on GuacaMol. LSTM, NAGVAE, and MCTS are tailored for molecule datasets; ConGress, DiGress, and DisCo are general graph generation models.
@@
DisCo-MPNN 68.7 100.0 96.4 77.0 36.4
Results and analysis.
Table 2 shows the performance on QM9 dataset. Our observation is consistent with the performance comparison on plain datasets: (1) DisCo-GT obtains slightly better or at least competitive performance against DiGress due to the shared graph-to-graph backbone, but our framework offers extra flexibility in the sampling process; (2) DisCo-MPNN obtains decent performance in terms of Validity, Uniqueness, and Novelty comparing with DisCo-GT.
Tables 3 and 4 show the performance on MOSES and GuacaMol datasets which further verifies that the proposed DiSCo’s performance is on par with the SOTA general graph generative models, DiGress.
@@ 4.3 Efficiency Study
A major computation bottleneck is the graph-to-graph backbone $p_{0|t}^{\theta}$, which is based on graph transformer (GT) or message-passing neural networks (MPNN). We compare the number of parameters, the inference time, and the back-propagation time of GT and MPNN in Table 5. For a fair comparison, we set all the hidden dimensions of GT and MPNN as $256$ and the number of layers of both models as $5$. We use the Community [68] dataset and set the batch size as $64$. Table 5 shows that GT has a larger model capacity and more parameters at the expense of more expensive model training and inference.
Table 5: Efficiency comparison of backbone models on community dataset. Inference and Backpropagation are the time (second/iteration) for model inference and backpropagation training.
@@
# of Parameters 14,216,707 7,047,427
Inference 0.065 0.022
Backpropagation 0.034 0.018
Table 6: Ablation study (%) on QM9 dataset. V., U., and N. mean Valid, Unique, and Novel.
Backbone Ref. Dist. Sampling steps Valid $\uparrow$ V.U. $\uparrow$ V.U.N. $\uparrow$
GT Marginal 500 99.3$\pm$0.6 98.9$\pm$0.6 56.2$\pm$0.4
@@ 4.4 Ablation Study
An ablation study on DisCo with different backbones (GT vs. MPNN), reference distributions (marginal vs. uniform), and the number of sampling steps ($1$, $5$, $10$, $30$, $100$, and $500$) is presented in Table 6. The number of sampling steps is determined by $\tau$, which is $\texttt{round}(\frac{1}{\tau})$ if the converged time is set as $T=1$. Overall, we observe the following results.
  • GT overall performs better than MPNN as the backbone model in terms of effectiveness.
  • Generally, the fewer sampling steps, the lower the generation quality. In some cases (e.g., the marginal reference distribution) with the sampling steps varying significantly (e.g., from $500$ to $30$), the performance change is still very slight, implying our method’s high robustness in sampling steps.
  • The marginal reference distribution is better than the uniform distribution, consistent with the conclusion from DiGress [62].
@@ 5 Related Work
We present a brief literature review on diffusion generative models and their applications for graph generation tasks.
Diffusion models are a family of popular generative models [66]. Their core idea is progressively adding noise into the data and designing a parametric model to sample clean data from tractable noise reversely. Its formulation can be either interpreted from the score-matching [58, 59] or the variational autoencoder perspective [18]. Beyond the continuous-state space where the diffusion process corrupts the given data into Gaussian noise [57], some efforts propose to design the diffusion process in the discrete-state space [19]. For example, D3PM [2] designs the discrete diffusion process by cumulative product of a fixed transition probability matrix; $\tau$-LDR [7] generalizes D3PM by formulating the diffusion process as a continuous-time Markov chain.
Diffusion generative models attract broad attention for graph generation tasks [37], for their promising applications such as molecule design [55, 20, 22, 36]. Pioneering works such as EDP-GNN [48] and GDSS [27] diffuse graph data in a continuous state space [21]. DiGress [62] is the first one-shot diffusion model to formulate graph data in discrete state space, followed by a very recent work MCD [39], both in the discrete-time setting. DruM [26] proposes to mix the diffusion process by conditioning on the endpoint in the data distribution. Edge [9] designs the diffusion process by diffusing graphs into empty graphs. Interestingly, GRAPHARM [29] uses an autoregressive diffusion model for graph generation. Besides, Liu et al. [38] apply the diffusion models for molecule property prediction. In addition to these works, there are many other task-tailored graph diffusion generative models [40, 25, 34, 51, 65, 64, 3, 63]. Interested readers are referred to this survey [37].
@@ 6 Conclusion
This paper introduces the first discrete-state continuous-time graph diffusion generative model, DisCo. Our model effectively marries continuous-time Markov Chain formulation with the discrete nature of graph data, addressing the fundamental sampling limitation of prior models. DisCo’s training objective is concise with a solid theoretical foundation. We also propose a simplified message-passing architecture to serve as the graph-to-graph backbone, which theoretically has desirable properties against permutation of node ordering and empirically demonstrates competitive performance against existing graph generative models in tests on various datasets.
7 Broader Impact
@@
  • Section A details the derivation of the factorized rate matrices.
  • Section B details the forward transition probability matrix given a rate matrix.
  • Section C includes all the proofs.
@@
  • Section D includes a step-by-step sampling algorithm.
  • Section E details the auxiliary features and neural modules used by DisCo.
  • Section F includes detailed experimental settings and additional experimental results.
@@
    • Section F.4: additional effectiveness evaluation
@@
  • Section G discusses this paper’s limitations and future work.
@@ Appendix C Proofs
$=\mathbf{I}+\Delta t\beta(\xi)\mathbf{R}_{f}+O((\Delta t)^{2}),$ (18)
where the second line we use the Mean Value Theorem. If the high-order term $O((\Delta t)^{2})$ is omitted and we short $\beta_{\Delta t}=\Delta t\beta(\xi)$, for $\mathbf{R}_{f}=\mathbf{1}\mathbf{1}^{\top}-b\mathbf{I}$, we have
@@
Because $\sum_{j=1}^{n}\mathbf{r}_{j,i}=\sum_{j=1}^{n}\mathbf{r}_{\sigma(j),\sigma(i)}$, $\mathbf{r}_{i}\odot\mathbf{r}_{j}=\mathbf{r}_{\sigma(i)}\odot\mathbf{r}_{\sigma(j)}$, and the FiLM module (Eq. (75)) is not related to the node ordering,
@@
where (*) is based on the permutation invariant of the forward process and its rate matrix and (**) is based on the permutation equivariance of the graph-to-graph backbone $p^{\theta}_{0|t}$. ∎
@@
We follow the notation and present the permutation $\mathcal{P}$ on every node as a bijection $\sigma:\{1,\dots,n\}\mapsto\{1,\dots,n\}$. We first analyze the cross-entropy loss on the nodes for a single training graph $\mathcal{G}_{0}$
@@ Appendix D Sampling Algorithm
7:              $J_{f^{i},s}\leftarrow\textrm{Poisson}(\tau\mathbf{R}_{t}^{i}(s,f^{i}))$triangleright # of transition for every edge
@@
Appendix E Auxiliary Features and Implementation Details
@@ F.1 Hardware and Software
We implement DisCo in PyTorch [^2] https://pytorch.org/ and PyTorch-geometric [^3] https://pytorch-geometric.readthedocs.io/en/latest/. All the efficiency study results are from one NVIDIA Tesla V100 SXM2-32GB GPU on a server with 96 Intel(R) Xeon(R) Gold 6240R CPU @ 2.40GHz processors and 1.5T RAM.
@@ F.2.2 Detailed Settings on Plain Graph Datasets
where ($(\{\mathcal{G}\}_{\texttt{gen}}$, ($(\{\mathcal{G}\}_{\texttt{train}}$, and ($(\{\mathcal{G}\}_{\texttt{test}}$ are the sets of generated graphs, training graphs, and test graphs, respectively. We report the above relative squared MMD for degree distributions (Deg.), clustering coefficient distributions (Clus.), and average orbit counts (Orb.) statistics (the number of occurrences of all substructures with $4$ nodes), which are standard metrics [68] for unattributed graphs in Table 1.
In addition, the Uniqueness, Novelty, and Validity are chosen. Uniqueness reports the fraction of the generated nonisomorphic graphs; Novelty reports the fraction of the generated graphs not isomorphic with any graph from the training set; Validity checks the fraction of the generated graphs following some specific rules. For the SBM dataset, we follow the validity check from [43] whose core idea is to check whether real SBM graphs are statistically indistinguishable from the generated graphs; for the Planar dataset, we check whether the generated graphs are connected and are indeed planar graphs. Because the Community dataset does not have the Validity metric, we only report the Uniqueness, Novelty, and Validity results on the SBM and Planar datasets.
@@ F.3 Hyperparameter Settings
As we introduced in Proposition 3.2, we tried two sets of rate matrices for the node and edge forward diffusion, so that the converged distribution is either uniform or marginal distribution. We found the marginal distribution leads to better results than the uniform distribution. Thus, the reference distribution is the marginal distribution for all the main results, except Table 6. The performance comparison between the marginal diffusion and uniform diffusion is presented in the Ablation study in Section  4.4. The $\beta(t)$ controls how fast the forward process converges to the reference distribution, which is set as $\beta(t)=\alpha\gamma^{t}log(\gamma)$, which is consistent with many existing works [18, 59, 7]. In our implementation, we assume the converged time $T=1$ and for the forward diffusion hyperparameters $(\alpha,\gamma)$ we tried two sets: $(1.0,5.0)$ and $(0.8,2.0)$ where the former one can ensure at $T=1$ the distribution is very close to the reference distribution, and the latter one does not fully corrupt the raw data distribution so the graph-to-graph model $p_{0|t}^{\theta}$ is easier to train.
@@
For DisCo-GT, the parametric graph-to-graph model $p^{\theta}_{0|t}$ is graph transformer (GT)-based. We use the exactly same GT architecture as DiGress [62] and adopt their recommended configurations  [^7] https://github.com/cvignac/DiGress/tree/main/configs/experiment. The reason is that this architecture is not our contribution, and setting the graph-to-graph model $p^{\theta}_{0|t}$ same can ensure a fair comparison between the discrete-time graph diffusion framework (from DiGress) and the continuous-time graph diffusion framework (from this work). For DisCo-MPNN, we search the number of MPNN layers from $\{3,5,8\}$, set all the hidden dimensions the same, and search it from $\{256,512\}$. For both variants, the dropout is set as $0.1$, the learning rate is set as $2e^{-4}$, and the weight decay is set as 00.
Table 8: Generation performance on the Community dataset.
@@ F.4 Additional Results on Community
Additional Community plain graph dataset results are in Table 8. Our observation is consistent with the main content: both variants of DisCo are on par with the SOTA general graph diffusion generative model, DiGress.
F.5 Convergence Study

### v2

@@ 4.2 Molecule Graph Generation
Table 4: Performance on GuacaMol. LSTM, NAGVAE, and MCTS are tailored for molecule datasets; ConGress, DiGress, and DisCo are general graph generation models.
@@
Results. Table 2 shows the performance on QM9 dataset. Our observation is consistent with the performance comparison on plain datasets: (1) DisCo-GT obtains slightly better or at least competitive performance against DiGress due to the shared graph-to-graph backbone, but our framework offers extra flexibility in the sampling process; (2) DisCo-MPNN obtains decent performance in terms of Validity, Uniqueness, and Novelty comparing with DisCo-GT.
Tables 3 and 4 show the performance on MOSES and GuacaMol which further verifies that (1) performance of DisCo-GT is on par with the SOTA general graph generative models, DiGress and (2) DisCo-MPNN has decent performance, but worse than DisCo-GT and DiGress.
@@ 4.3 Efficiency Study
Table 5: Efficiency comparison in terms of number of parameters, forward and backpropagation time (second/iteration).
@@
# Parameters $14\times 10^{6}$ $7\times 10^{6}$
Forward $0.065$ $0.022$
Backprop. $0.034$ $0.018$
A major computation bottleneck is the graph-to-graph backbone $p_{0|t}^{\theta}$, which is GT or MPNN. We compare the number of parameters, the forward and back-propagation time of GT and MPNN in Table 5. For a fair comparison, we set all the hidden dimensions of GT and MPNN as $256$ and the number of layers as $5$. We use the Community [82] dataset and set the batch size as $64$. Table 5 shows that GT has a larger capacity and more parameters at the expense of more expensive training.
Table 6: Ablation study (mean±$\pm$±std%) with GT backbone. V., U., and N. mean Valid, Unique, and Novel.
Ref. Dist. Steps Valid $\uparrow$ V.U. $\uparrow$ V.U.N. $\uparrow$
Marginal 500 99.3$\pm$0.6 98.9$\pm$0.6 56.2$\pm$0.4
@@ 4.4 Ablation Study
An ablation study on DisCo-GT for reference distributions (marginal vs. uniform), and sampling steps ($1$ to $500$) is presented in Table 6. The number of sampling steps is $\texttt{round}(\frac{1}{\tau})$ if $T=1$. QM9 dataset is chosen. A similar ablation study on DisCo-MPNN is in Table 9 in Appendix. We observe that first, generally, the fewer sampling steps, the lower the generation quality. In some cases (e.g., the marginal distribution) with the sampling steps decreasing significantly (e.g., from $500$ to $30$), the performance degradation is still very slight, implying our method’s high robustness in sampling steps. Second, the marginal reference distribution is better than the uniform distribution, consistent with the observation from DiGress [73].
@@ 5 Related Work
Diffusion models [80] can be interpreted from both the score-matching [69, 70] or the variational autoencoder perspective [23, 35, 34]. Pioneering efforts on diffusion generative modeling study the process in continuous-state [67, 23, 68] whose typical reference distribution is Gaussian. Beyond that, some efforts propose discrete-state models [24] to . E.g., D3PM [2] designs the discrete diffusion process by multiplication of transition matrices; $\tau$-LDR [8] generalizes D3PM by formulating a continuous-time Markov chain;  [71] proposes a singleton conditional distribution-based objective for the continuous-time Markov chain-based model whose rationale is Brook’s Lemma [5, 49].
Diffusion models are widely used in graph generation tasks [44, 13, 85, 86, 84, 15, 83, 14, 61, 74, 47, 79, 78, 59, 3] such as molecule design [65, 25, 27, 43]. Pioneering works such as EDP-GNN [56] and GDSS [32] diffuse graph data in a continuous state space [26]. DiscDDPM [22] is an early effort to modify the DDPM architecture into a discrete state. In addition, DiGress [73] is also a one-shot discrete-state diffusion model, followed by a very recent work MCD [46], both in the discrete-time setting. Beyond the above-mentioned efforts, DruM [31] proposes to mix the diffusion process. EDGE [10] proposes an interesting process: diffusing graphs into empty graphs. Besides, GRAPHARM [36] proposes an autoregressive graph diffusion model, and  [45] applies the diffusion models for molecule property prediction tasks. In addition to the above-mentioned general graph diffusion models, there are many other task-tailored graph diffusion generative models [48, 30, 41, 60, 77, 76, 4, 75], which incorporate more in-depth domain expertise into the model design. Interested readers are referred to this survey [44].
@@ 6 Conclusion
This paper introduces the first discrete-state continuous-time graph diffusion generative model, DisCo. Our model effectively marries continuous-time Markov Chain formulation with the discrete nature of graph data, addressing the fundamental sampling limitation of prior models. DisCo’s training objective is concise with a solid theoretical foundation. We also propose a simplified message-passing architecture to serve as the graph-to-graph backbone, which theoretically has desirable properties against permutation of node ordering and empirically demonstrates decent performance against existing graph generative models in tests on various datasets.
Acknowledgments and Disclosure of Funding
ZX, RQ, ZZ, and HT are partially supported by NSF (2324770). The content of the information in this document does not necessarily reflect the position or the policy of the Government, and no official endorsement should be inferred. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation here on.
@@
  • Section A: the derivation of the factorized rate matrices.
  • Section B: the forward transition probability matrix given a rate matrix.
  • Section C: all the proofs.
@@
  • Section D: a step-by-step sampling algorithm.
  • Section E: the auxiliary features and neural modules used by DisCo.
  • Section F: detailed experimental settings and additional experimental results.
@@
    • Section F.4: additional effectiveness evaluation on Community Dataset
    • Section F.5: additional ablation study with the MPNN backbone
@@
  • Section G: this paper’s limitations and future work.
  • Section H: the broad impact of this paper.
@@ Appendix C Proofs
=$\overset{(*)}{=}\mathbf{I}+\Delta t\beta(\xi)\mathbf{R}_{f}+O((\Delta t)^{2}),$ (14)
where (*) is based on the Mean Value Theorem. If the high-order term $O((\Delta t)^{2})$ is omitted and we short $\beta_{\Delta t}=\Delta t\beta(\xi)$, for $\mathbf{R}_{f}=\mathbf{1}\mathbf{1}^{\top}-b\mathbf{I}$, we have
@@
Because $\sum_{j=1}^{n}\mathbf{r}^{(j,i)}=\sum_{j=1}^{n}\mathbf{r}^{(\sigma(j),\sigma(i))}$, $\mathbf{r}^{i}\odot\mathbf{r}^{j}=\mathbf{r}^{\sigma(i)}\odot\mathbf{r}^{\sigma(j)}$, and the FiLM module (Eq. (71)) is not related to the node ordering,
@@
where (*) is based on the permutation invariant of the forward process and its rate matrix; (**) is based on the permutation equivariance of the graph-to-graph backbone $p^{\theta}_{0|t}$. ∎
@@
We follow the notation and present the permutation $\mathcal{P}$ on every node as a bijection $\sigma:\{1,\dots,n\}\mapsto\{1,\dots,n\}$. We first analyze the cross-entropy loss on the nodes for a single training graph $\mathcal{G}_{0}$ and taking expectation $\mathbb{E}_{\mathcal{G}_{0}}$ keeps the permutation invariance:
@@ Appendix D Sampling Algorithm
7:              $J_{f^{i},s}\leftarrow\textrm{Poisson}(\tau\mathbf{R}_{t}^{i}(s,f^{i}))$triangleright # of transition for every node
@@
Appendix E Auxiliary Features, PNA and FiLM Modules
@@
Next, two important modules used in the MPNN backbone: PNA and FiLM are detailed.
@@ F.1 Hardware and Software
We implement DisCo in PyTorch [^3] https://pytorch.org and PyTorch-geometric [^4] https://pytorch-geometric.readthedocs.io/en/latest. All the efficiency study results are from one NVIDIA Tesla V100 SXM2-32GB GPU on a server with 96 Intel(R) Xeon(R) Gold 6240R CPU @ 2.40GHz processors and 1.5T RAM. The training on QM9 and Community can be finished in 2 hours. For the training on SBM, Planar, it can be finished within 48 hours to get decent validity. The training on MOSES and GuacaMol can be finished within 96 hours.
@@ F.2.2 Detailed Settings on Plain Graph Datasets
where ($(\{\mathcal{G}\}_{\texttt{gen}}$, ($(\{\mathcal{G}\}_{\texttt{train}}$, and ($(\{\mathcal{G}\}_{\texttt{test}}$ are the sets of generated graphs, training graphs, and test graphs, respectively. We report the above relative squared MMD for degree distributions (Deg.), clustering coefficient distributions (Clus.), and average orbit counts (Orb.) statistics (the number of occurrences of all substructures with $4$ nodes). In addition, the Uniqueness, Novelty, and Validity are chosen. Uniqueness reports the fraction of the generated nonisomorphic graphs; Novelty reports the fraction of the generated graphs not isomorphic with any graph from the training set; Validity checks the fraction of the generated graphs following some specific rules. For the SBM dataset, we follow the validity check from [51] whose core idea is to check whether real SBM graphs are statistically indistinguishable from the generated graphs; for the Planar dataset, we check whether the generated graphs are connected and are indeed planar graphs. Because the Community dataset does not have the Validity metric, we only report the Uniqueness, Novelty, and Validity results on the SBM and Planar datasets.
We report mean±$\pm$±std in $5$ runs.
Baseline methods.
GraphRNN [82], GRAN [42], GG-GAN [37], MolGAN [9], SPECTRE [51], EDP-GNN [56], GraphGDP [26], DiscDDPM [22], EDGE [10], ConGress [73], DiGress [73] are chosen.
@@ F.2.3 Detailed Settings on Molecule Graph Datasets
We report mean±$\pm$±std in $5$ runs except MOSES and GuacaMol, whose computations are too expensive to repeat multiple times.
Baseline methods.
CharacterVAE [20], GrammarVAE [38], GraphVAE [66], GT-VAE [55], Set2GraphVAE [72], GG-GAN [37], MolGAN [9], SPECTRE [51], GraphNVP [50], GDSS [32], EDGE [10], ConGress [73], DiGress [73], GRAPHARM [36],VAE [21], JT-VAE [29], GraphINVENT [53], LSTM [64], NAGVAE [40], and MCTS [28] are chosen.
@@ F.3 Hyperparameter Settings
As we introduced in Proposition 3.2, we tried two sets of rate matrices for the node and edge forward diffusion, so that the converged distribution is either uniform or marginal distribution. We found the marginal distribution leads to better results than the uniform distribution. Thus, the reference distribution is the marginal distribution for all the main results, except Tables 6 and 9. The performance comparison between the marginal diffusion and uniform diffusion is presented in the ablation study in Sections  4.4 andF.5. The $\beta(t)$ controls how fast the forward process converges to the reference distribution, which is set as $\beta(t)=\alpha\gamma^{t}log(\gamma)$, which is consistent with many existing works [23, 70, 8]. In our implementation, we assume the converged time $T=1$ and for the forward diffusion hyperparameters $(\alpha,\gamma)$ we tried two sets: $(1.0,5.0)$ and $(0.8,2.0)$ where the former one can ensure at $T=1$ the distribution is very close to the reference distribution, and the latter one does not fully corrupt the raw data distribution so the graph-to-graph model $p_{0|t}^{\theta}$ is easier to train.
@@
For DisCo-GT, the parametric graph-to-graph model $p^{\theta}_{0|t}$ is graph transformer (GT). We use the exactly same GT architecture as DiGress [73] and adopt their recommended configurations  [^8] https://github.com/cvignac/DiGress/tree/main/configs/experiment. The reason is that this architecture is not our contribution, and setting the graph-to-graph model $p^{\theta}_{0|t}$ same can ensure a fair comparison between the discrete-time graph diffusion framework (from DiGress) and the continuous-time graph diffusion framework (from this work). For DisCo-MPNN, we search the number of MPNN layers from $\{3,5,8\}$, set all the hidden dimensions the same, and search it from $\{256,512\}$. For both variants, the dropout is set as $0.1$, the learning rate is set as $2e^{-4}$, and the weight decay is set as 00.
Table 8: Generation performance (mean±$\pm$±std) on the Community dataset.
@@ F.4 Additional Results on Community
Additional Community plain graph dataset results are in Table 8. Our observation is consistent with the main content: both variants of DisCo are on par with, or even better than the SOTA general graph diffusion generative model, DiGress.
F.5 Additional Ablation Study
Table 9 shows the ablation study of DisCo-MPNN on QM9 dataset. Our observations are consistent with the main content: (1) generally, the fewer sampling steps, the lower the generation quality but method’s performance is robust in terms of the decreasing of sampling steps; (2) the marginal reference distribution is better than the uniform distribution, consistent with the observation from DiGress [73].
Table 9: Ablation study (mean±$\pm$±std%) with MPNN backbone. V., U., and N. mean Valid, Unique, and Novel.
Ref. Dist. Steps Valid $\uparrow$ V.U. $\uparrow$ V.U.N. $\uparrow$
Marginal 500 98.9$\pm$0.7 98.7$\pm$0.5 68.7$\pm$0.2
100 98.4$\pm$1.1 98.0$\pm$1.0 69.1$\pm$0.6
30 97.7$\pm$1.2 97.5$\pm$0.8 70.4$\pm$1.1
10 92.3$\pm$1.9 91.9$\pm$2.2 66.4$\pm$1.7
5 88.8$\pm$3.3 87.1$\pm$2.8 67.3$\pm$2.9
1 64.4$\pm$2.7 63.2$\pm$1.9 55.8$\pm$1.4
Uniform 500 93.5$\pm$1.7 93.2$\pm$1.1 64.9$\pm$1.0
100 93.1$\pm$2.1 92.6$\pm$1.7 66.2$\pm$1.9
30 87.1$\pm$1.8 86.8$\pm$1.1 64.0$\pm$1.0
10 83.7$\pm$3.2 81.9$\pm$2.1 61.3$\pm$2.0
5 81.5$\pm$2.9 75.4$\pm$3.4 64.6$\pm$2.3
1 71.3$\pm$2.3 42.2$\pm$4.0 36.9$\pm$3.2
F.6 Convergence Study